Assuming that there is no government spending or trade, an economy’s aggregate

demand is given by its domestic consumption C and investment I:AD=C+I=c0 +c1Y+I

In the economy’s goods market equilibrium this equals its output: AD = Y. Solving for Y, this yields:

Y = [1/(1 - c1 )] (c0 + I)

Given this equation, which of the following statements is correct? Select one answer and provide explanations for each choice:

  1. a) The multiplier is given by 1 – c1.

  2. b) The boost in the economy’s output is the same, regardless of whether the aggregate demand shock comes from an increase in investment I or in autonomous consumptionc0.

  3. c) The larger the marginal propensity to consume (c1), the smaller the multiplier.

  4. d) If c1 = 1/3, then a £1 million increase in investment would result in a £2 million increase in output.
Answer #1

(1) False

Multiplier = 1 / (1 - c1)

(2) True

Any change in autonomous expenditure will cause the same change in output.

(3) False

The larger the value of c1, the smaller the value of (1 - c1) and so the larger the value of [1 / (1 - c1)] i.e. multiplier.

(4) False

Multiplier = 1 / [1 - (1/3)] = 1 / [(3 - 1) / 3] = 1 / (2/3) = 3/2 = 1.5

So, when I increases by 1 million, output increases by (1 x 1.5) = 1.5 million.